Part A. Considering light at the two ends of the visible light spectrum, violet light has...

Part A.

Considering light at the two ends of the visible light spectrum, violet light has a _____ wavelength and a _____ photon energy than red light.

  A.Shorter, higher
  B.Longer, higher
C. Longer, lower
D. Shorter, lower
   
   
   
 



Part B.

Lenz’s law states that an induced magnetic field in a conductor

A. Opposes the change in flux through the conductor.
B. Opposes the applied magnetic field.
C. Induces a matching electric field that causes a current.
D. Opposes the applied flux through the conductor.

Part C.

In this video, we see that a _____ magnetic field can create an electric current.

In this video, we see that a _____ magnetic field can create an electric current.

  A.Perpendicular
  B.Changing
  C.Dipole
  D.Strong

Part D.

A particle with a charge of q = -5.90 nC is moving in a uniform magnetic field of B⃗ =( -1.25 T ) k^.The magnetic force on the particle is measured to be F⃗ =( −7.60×10−7 N )j^.

Calculate vx, the x component of the velocity of the particle?
Answer #1

Part A

In the visible pattern, the wavelength increases from violet to Red.

The wavelength is inversely proportional to energy

violet light has a shorter wavelength and a higher photon energy than red light.

Correct option: A

-------------------------------------------------------------------------------------------------------------------------------------

Part B:

Lenz’s law states that an induced magnetic field in a conductor,

Opposes the change in flux through the conductor.

Correct option is ( A )

-------------------------------------------------------------------------------------------------------------------------------------------

Part C.

Vedio not appeared in the question to answer this question

------------------------------------------------------------------------------------------------------------------------------------------

Part D:

The magnetic force on the particle is

F = q ( v X B )

( −7.60×10−7 N )j^. = -5.90 ×10−9 ( v i x ( -1.25 T ) k^ )

Solve for v

v = 103 m /s